Qd= Qs given the following equations: 400 - 2P + 4j = - 100 + 3P and 400+400+100 = 3P+2P calculate price and quantity
1
Expert's answer
2021-08-24T10:16:26-0400

"400-2p+4j=-100\\400+4j+100=2p\\p=200+2j+50"


"400+400+100=3p+2p\\900=5p\\p=180"


"200+2j+50=180\\250+25=180\\2j=180-250\\2j=70\\j=35"